Matthias E. Johnson

matthias@opennomad.com https://opennomad.com/

Objective

I thrive working on interesting problems of scale and complexity. I desire to build platforms that matter and leverage modern technologies with a focus on simplicity and high degrees of automation.

Experience

Independent DevOps and Cloud consultant 10/2015 - present

• General Consulting

• Environment Deep Dives and Reviews • Performance and Stability improvements • CI/CD implementation

Site Reliability Engineering Architect - ReactiveOps 10/2015 - 12/2017

• Client Consulting

• Development

DevOps Manager/Lead/Architect (title/role change) - Catalyst Repository Systems 1/2012 - 10/2015

• Architecture and Design

• Development

• Automation

• Monitoring and Reporting

• Management ...

Search Engine Architect / Manager of Search Team - Catalyst Repository Systems 7/2007 - 8/2013

• Search Engine Architect

• Monitoring and Reporting

• Automation

• Capacity Planning

• Supervision and Management

Manager of Linux Enterprise Hosting (title/role change) - ViaWest, Inc. 3/2007 - 7/2007

• Linux Architect

• Product Development and Planning

• Project Management

• Tier 3 support

• Mentoring

Manager of Technical Operations (title/role change) - ViaWest, Inc. 6/2006 - 3/2007

• Management of Senior Technical Customer Support Team • Product Development and Planning

• Project Management

• Tier 3 support

Manager of Systems Administration (internal promotion) - ViaWest, Inc. 8/2005 - 6/2006

• Management of Systems Administration Group

• Product Development and Planning

• Project Management

• Tier 3 support

Systems Engineer - ViaWest, Inc. 6/2004 - 8/2005

• Senior Lead for Infrastructure Services

• Member of Hosting Group

• Strategic Research

• E-mail administrator

• Tier 3 support Prior work experience has been removed. For complete employment history please see the full version of my resume.

Technical Skills

I have experience with a tremendous number of systems, services, and languages. The following list is heavily abbreviated and reflects my current focus. The full version of my resume details many of these in the expanded experience section. Programming Languages Ruby, JavaScript, Bash, Perl, Python System Platforms Ubuntu, Amazon Linux, RHEL, Windows, Solaris, True64 Technology Stacks AWS, Docker, Kubernetes, Terraform, Ansible, ELK Services Elasticsearch, Redis, RabbitMQ, Logstash, many RDBMS

Education

MS in Computer Science Minnesota State University, Mankato - 5/2000 Completed thesis entitled Knowledge extraction from Neural Nets through Graph Theory.

BS in Computer Science Minnesota State University, Mankato - 3/1997 Received grant for research on Convex Subsets as a heuristic for Tournament Isomorphism presented as poster session at the 1996 Small College Computing Symposium in St. Cloud, MN. Another research project, related to Data Mining, on Explanation Facilities in Neural nets guided further research at Minnesota State University.